OpenStudy (mpj4):

How to solve 2 equations with 4 unknowns?

OpenStudy (mpj4):

\[3t - 4u + v = 2\] \[t + u - 2v + 3w = 1\]

geerky42 (geerky42):

What do you need to do? I don't think solving each variable is possible.

OpenStudy (cwrw238):

yes I agree

OpenStudy (mpj4):

It has an infinite set of solutions. And there's an answer key on the book.

OpenStudy (mpj4):

\[( c_1 - (12/7) c_2 + (6/7), c_1 - (9/7)c_2 + (1/7), c_1, c_2) \]

geerky42 (geerky42):

What is full question?

OpenStudy (mpj4):

solve each system of equations by the Gaussian elimination method.

OpenStudy (mpj4):

It's in the topic of matrices so I tried solving it using an augmented matrix. But it was too hard.

ganeshie8 (ganeshie8):

which part is hard ?

OpenStudy (mpj4):

The part where I try to make sense of it all

ganeshie8 (ganeshie8):

say you're given only 1 equation wih 4 unknowns to solve would it make sense to you ?

OpenStudy (mpj4):

2 equations actually

OpenStudy (mpj4):

Ah, I give up. It'll be better to use my time to study other things xD

ganeshie8 (ganeshie8):

\[\left[\begin{matrix} 3&-4&1&0&|&2\\ 1&1&-2&3&|&1 \end{matrix}\right] \]

ganeshie8 (ganeshie8):

is that your augmented matrix ?

OpenStudy (mpj4):

yes

ganeshie8 (ganeshie8):

lets row reduce and find `particular` and `null` solutions heard of them before ?

OpenStudy (mpj4):

I've only been given the basics so, nope.

ganeshie8 (ganeshie8):

\[\left[\begin{matrix} 3&-4&1&0&|&2\\ 1&1&-2&3&|&1 \end{matrix}\right] \] R2-(1/3)R1 : \[\left[\begin{matrix} 3&-4&1&0&|&2\\ 0&7/3&-7/3&3&|&1/3 \end{matrix}\right] \]

ganeshie8 (ganeshie8):

its easy, notice that you have two pivots. so there will be two free variables in the general solution

OpenStudy (mpj4):

Ahh, if that's what you're talking about then, I know.

ganeshie8 (ganeshie8):

yes you just need to know finding null and particular solutions, thats all

ganeshie8 (ganeshie8):

lets find the null solution first

ganeshie8 (ganeshie8):

you want to solve below for null solution : \[ \left[\begin{matrix} 3&-4&1&0\\ 0&7/3&-7/3&3 \end{matrix}\right] \begin{pmatrix} t\\ u\\ v\\ w \end{pmatrix} = \left[\begin{matrix} 0\\ 0 \end{matrix}\right] \]

ganeshie8 (ganeshie8):

know how to find the null solution ?

OpenStudy (mpj4):

I'm searching to see if I do.

ganeshie8 (ganeshie8):

let me tell you the strategy quick : 1) set v=0, w=1 and find one solution 2) set v=1, w=0 and find another solution the combination of above solutions is the complete null solution

ganeshie8 (ganeshie8):

v = 0, w = 1 : \[ \left[\begin{matrix} 3&-4&1&0\\ 0&7/3&-7/3&3 \end{matrix}\right] \begin{pmatrix} t\\ u\\ 0\\ 1 \end{pmatrix} = \left[\begin{matrix} 0\\ 0 \end{matrix}\right] \]

ganeshie8 (ganeshie8):

can you solve t and u ?

OpenStudy (mpj4):

Ok, gimme some time if it's ok

ganeshie8 (ganeshie8):

Okay ill show you how to solve one solution, maybe you can try the other solution

ganeshie8 (ganeshie8):

v = 0, w = 1 : \[ \left[\begin{matrix} 3&-4&1&0\\ 0&7/3&-7/3&3 \end{matrix}\right] \begin{pmatrix} t\\ u\\ 0\\ 1 \end{pmatrix} = \left[\begin{matrix} 0\\ 0 \end{matrix}\right] \] 3t - 4u = 0 7/3u +3 = 0 solving you get : u = -9/7 t = -12/7

ganeshie8 (ganeshie8):

so one null solution vector is : \[\begin{pmatrix} -12/7\\ -9/7\\ 0\\ 1\\ \end{pmatrix}\]

ganeshie8 (ganeshie8):

see if you can find the other null solution by letting v=1 and w=0, you need to solve below equation : \[\left[\begin{matrix} 3&-4&1&0\\ 0&7/3&-7/3&3 \end{matrix}\right] \begin{pmatrix} t\\ u\\ 1\\ 0 \end{pmatrix} = \left[\begin{matrix} 0\\ 0 \end{matrix}\right]\]

OpenStudy (mpj4):

3t -4u +1 = 0 7/3 u - 7/3 = 0 u = 1 t = 1

ganeshie8 (ganeshie8):

yes so the second null solution vector is \[\begin{pmatrix} 1\\ 1\\ 1\\ 0\\ \end{pmatrix}\]

ganeshie8 (ganeshie8):

Can we say the complete `nullspace` is : \[ c_1\begin{pmatrix} 1\\ 1\\ 1\\ 0\\ \end{pmatrix} + c_2 \begin{pmatrix} -12/7\\ -9/7\\ 0\\ 1\\ \end{pmatrix} \] ?

OpenStudy (mpj4):

\[\left(\begin{matrix}-12/7 \\ -9/7\\ 1\\ 1 \end{matrix}\right)\]?

ganeshie8 (ganeshie8):

no don't combine, null solution is a linear combination of the two solution vectors we have found earlier

ganeshie8 (ganeshie8):

you need to find a particular solution and add it to null solution for the final general solution

ganeshie8 (ganeshie8):

lets find the particular solution quick

ganeshie8 (ganeshie8):

for particular solution, you swt all the free variables to 0 and solve the original given equation

ganeshie8 (ganeshie8):

solve below for particular solution : \[ \left[\begin{matrix} 3&-4&1&0\\ 0&7/3&-7/3&3 \end{matrix}\right] \begin{pmatrix} t\\ u\\ 0\\ 0 \end{pmatrix} = \left[\begin{matrix} 2\\ 1/3 \end{matrix}\right] \]

OpenStudy (mpj4):

3t -4u = 2 7/3 u = 1/3 u - 1/7 t = 6/7 I really have no idea.

ganeshie8 (ganeshie8):

thats right! so a particular solution is \[ \begin{pmatrix} 6/7\\ 1/7\\ 0\\ 0\\ \end{pmatrix} \]

ganeshie8 (ganeshie8):

we're done! we have `particular` and `null` solutions, so we can write the final `general` solution

ganeshie8 (ganeshie8):

`general` = `particular` + `null`

ganeshie8 (ganeshie8):

general solution : \[ \begin{pmatrix} 6/7\\ 1/7\\ 0\\ 0\\ \end{pmatrix} + c_1\begin{pmatrix} 1\\ 1\\ 1\\ 0\\ \end{pmatrix} + c_2 \begin{pmatrix} -12/7\\ -9/7\\ 0\\ 1\\ \end{pmatrix} \]

ganeshie8 (ganeshie8):

see if that makes more or less sense..

OpenStudy (mpj4):

It made more sense than what I could have made. at least.

OpenStudy (mpj4):

Thanks a lot for the help.

ganeshie8 (ganeshie8):

basically it all boils down to just finding `particular` and `null` solutions

OpenStudy (mpj4):

Thanks, I have a test on it tomorrow anyway.

ganeshie8 (ganeshie8):

good luck!